A project was originally estimated to be completed at a cost of $325,000. The cost of what’s been completed so far is $250,000, and the project team has completed roughly 12 of the 16‐ week completion estimate.
Answer #1

250000/325000 = 77% of Budget at completion

The project has exhausted 71% of the total budget

12/16 = 75% of budgeted time

The project has completed 75% of the estimated time

earned value = 75% * 325000 = 243750

Cost variance = Earned value - actual cost = 243750- 2750000 = - $6250

The project is running behind budget and incurring more cost than projected/expected